# Enhancing Precision of Signal Correction in PVES Experiments: The Impact of Bayesian Analysis on the Results of the QWeak and MOLLER Experiments

### v1

@@ I Introduction
$A_{PV}=R_{\rm tot}\frac{\frac{A_{\rm expt}}{P_{b}}-\sum_{i}f_{i}^{\rm bkgd}A_{i}^{\rm bkgd}}{1-\sum_{i}f_{i}^{\rm bkgd}},$ (1)
where $A_{\rm expt}$ is the experimentally measured asymmetry, which must be corrected for background processes, characterized by fractional dilution factors, $f_{i}^{\rm bkgd}$, and asymmetries, $A_{i}^{\rm bkgd}$, $P_{b}$ is the beam polarization, and $R_{\rm tot}$ is an overall normalization factor.
@@ II.1 Bayesian Analysis
Bayesian analysis begins with Bayes’ theorem, which connects conditional probabilities, likelihood, priors, and the resulting posterior distributions. For model parameters $\theta$ and data $y$, Bayes’ theorem is formulated as:
$P(\theta|y)=\frac{P(y|\theta)P(\theta)}{P(y)},$ (2)
@@
where $P(\theta|y)$ is the posterior distribution of the model parameters after observing the data, $P(y|\theta)$ is the likelihood of the data given the parameters, $P(\theta)$ is the prior distribution, reflecting prior beliefs about the parameters, and $P(y)$ is a normalizing constant ensuring the posterior sums to one. This theorem demonstrates how beliefs are updated in light of new data and details how uncertainties regarding model parameters are quantified.
@@
To implement Bayesian analysis in the QWeak and MOLLER experiments, we quantify the prior distribution based on existing knowledge, use experimental data to form the likelihood function, and select a model to link these to the parameters of interest, the asymmetry components. Regarding the prior distribution, no prior knowledge is assumed for either experiment, allowing the measurements to be the dominant influence on the posterior distributions. Concerning the input data, the measured asymmetry values are crucial. These values depend on the configuration of the experimental setup and the number of detectors, which are detailed in subsequent sections. Alongside the measured asymmetry values, other parameters, such as the beam polarization angle and spin variation, should also be considered as input data. Finally, the models that link the input data to the model parameters differ between the experimental setups. The inherent complexities and computational challenges of working with these high-dimensional models necessitate the use of approximation techniques. These techniques aim to simplify reality, thereby enhancing our understanding of the system’s components. Various computational challenges in Bayesian inference have historically favoured frequentist methods until advancements in computational techniques allowed more efficient Bayesian methods. Techniques such as Approximate Bayesian Computation (ABC) [12], Integrated Nested Laplace Approximations (INLA) [13], and Variational Bayesian methods [14] have been developed. These methods generally approximate only marginal posterior distributions of individual parameters, not the joint distributions of multiple parameters with their potential correlations. MCMC methods, introduced by Gelfand and Smith in 1990 [15], have become prominent for their ability to approximate the joint posterior distribution through sampling. In these methods, which are computationally efficient and widely used in Bayesian analysis, the Monte Carlo part denotes the sampling process, and the Markov Chain part describes the mechanism for obtaining these samples.
@@ II.2 QWeak Experiment
In a shorter ancillary measurement of the QWeak experiment, the beam energy was increased to 3.35 GeV to measure the parity-violating asymmetry of inelastically scattered electrons (above the resonance region). This measurement had the aim of probing $\gamma-Z$ box diagram contributions to the asymmetry. In the interest of brevity, these inelastic data will be referred to in the following simply as “QWeak data”. At these kinematics, unlike for the primary weak charge measurement, pion production was not negligible, and so one of the 8 main detector modules was temporarily reconfigured as a pion detector, to allow the determination and correction for the asymmetry of the (approximately 0.1%) contamination from pions and their decay muons. A 10.2 cm thick lead wall was placed in front of the lowermost detector (MD7) to attenuate electrons more than pions because of their differing interactions with dense material, thereby enhancing the pion signal detected by MD7. Consequently, MD7 provided a unique response to the electron and pion mix that allows distinguishing between asymmetries caused by scattered electrons and pions.
@@
Here, $f_{\pi}^{i}$ represents the fractional yield of pions at MD $i$, and $A_{e(\pi)}^{L}$ and $A_{e(\pi)}^{T}$ denote the longitudinal and transverse asymmetries for electrons and pions, respectively. $\theta_{P}^{j}$ is the beam polarization angle for the run type $j$, and $f_{\text{NB}}^{i}$ is the neutral background yield fraction for MD $i$. The azimuthal angles, $\phi^{i}$, define the placement of the MDs with specific values like $\phi^{1}=0^{\circ}$ and $\phi^{2}=45^{\circ}$, among others. This model will be used to reanalyze the QWeak inelastic data, incorporating all inputs from Ref. [19].
@@ II.3 MOLLER Experiment
The experimental design differs significantly between the two; the QWeak data employed eight main detectors, with one acting as a pion detector, whereas the MOLLER experiment is planned to use a larger array of 84 main detector modules for ring 5 and 28 pion detector modules, necessitating an expanded Bayesian analysis framework. Note that in this paper, the analysis will focus on the main detector ring 5 and the pion detector. Eventually, we plan to expanded this analysis to include all six rings of the main detectors, comprising 224 modules, the pion detector, and the showermax, each including 28 modules.
@@ III Validation of the Results
This section illustrates the results of implementing Bayesian analysis using the inputs and models from the QWeak and MOLLER experiments. For QWeak, validation is achieved by comparing these results with those obtained from the (frequentist) Many-Worlds Monte Carlo minimization approach, as detailed in [19, 20]. For MOLLER, validation is conducted by testing agreement with the input asymmetries in the simulation.
@@ III.1 Reanalysis of the QWeak Inelastic Experiment
Incorporating all inputs from the QWeak inelastic data, $A^{\text{meas}}$, $f_{\pi}$, $f_{\text{NB}}$, $\theta$, $\phi$, and the model as described by Eq. 4, the asymmetry values associated with Møller electrons and pions, $A_{e}^{L}$, $A_{e}^{T}$, $A_{\pi}^{L}$, $A_{\pi}^{T}$, are calculated using two analytical methods: the Many-Worlds Monte Carlo Minimization method and Bayesian Analysis.
@@
Figure 1: Comparative analysis of measured and fitted asymmetry values across detectors for two data sets. On the left, the mixed data set is shown with measured asymmetry depicted as blue squares, compared with values derived through Bayesian inference (red circles) and Monte Carlo simulations (black triangles). On the right, the transverse data set is similarly represented. The data spans eight detectors in each plot, with error bars indicating the uncertainty for each data point. MD7 is highlighted with hollow symbols to indicate its distinct characteristics from the other MDs. Dashed lines in both data sets show the fitted sine wave in the Bayesian analysis, which is different from those obtained using the Monte Carlo method [19].
@@
This is explained by the mechanics of Bayesian analysis. As mentioned in the methodology, Bayes’ theorem requires an iterative process where the posterior distribution from each iteration is used as the prior for the next, incorporating new data sets. This method ensures that in the absence of initial prior knowledge, the Bayesian approach remains data-driven. Additionally, the uncertainty is quantified by the width of the posterior distribution in the final step, which is shaped by the data and results in more precise estimations. In contrast, the Monte Carlo minimization method derives component asymmetries from measured data by randomly sampling input values from Gaussian distributions based on their means and uncertainties. These values are used to calculate asymmetries, and the method minimizes the squared deviations between measured and calculated asymmetries to estimate the unknown components. This process produces distributions for each component. The RMS of these distributions is used to determine the uncertainty for each component, resulting in larger uncertainty estimates due to the broader spread of possible values.
@@ III.2 Analysis of the MOLLER Experiment
Incorporating the simulations and assumptions from the MOLLER experiment, the input data include $A^{\text{mock}}$, $f^{i}_{\pi}$, $\theta_{P}^{j}$, $\phi_{P}^{j}$, $A_{ij}$, $B_{ij}$, $C_{ij}$, and $D_{ij}$. The parameters to be extracted are the asymmetry values $A_{e}^{L}$, $A_{e}^{T}$, $A_{\pi}^{L}$, and $A_{\pi}^{T}$. The model, specified by Eq. 6, is applied separately to the pion and main detectors for both longitudinal and transverse measurements. Consequently, the asymmetry values are extracted using four equations: two for the pion detector and two for the main detector, covering both longitudinal and transverse measurements. The results of the extracted asymmetry values and their associated uncertainties in ppb are summarized in Table 2 and compared with the inputs of the analysis.
@@
Another way to verify the results is by substituting the asymmetry values along with other parameters into the right-hand side of Eq. 6, computing and deriving another set of asymmetry values, referred to as the fitted asymmetry values, similar to the case of Qweak. Figures 2 and 3 provide a comparative analysis of fitted versus mock asymmetry for the main and pion detectors in both the longitudinal and transverse data sets. In the top plots, the blue squares represent the average of mock asymmetry values, while the red circles denote the average of fitted asymmetry values obtained through Bayesian analysis. Due to the large scale of the asymmetry values, the error bars are too small to be visible in some of these plots. To address this, the middle plots subtract the mock asymmetry average from both the mock and fitted asymmetry values, effectively re-centering the data around zero. This re-centering provides a clearer visualization of the error bars, illustrating the measurement precision. In these plots, the blue dots represent the residual mock asymmetry values, which are zero as they are subtracted from themselves. The red circles display the residual fitted asymmetry values, with the error bars indicating the uncertainties. The bottom plots present histograms of the normalized residuals, defined as the difference between the fitted and mock asymmetry values, normalized by the uncertainty in the mock asymmetry values. These histograms help assess the normality and spread of these residuals.
@@ IV Discussion
The Bayesian approach showed more precise estimates and reduced uncertainties, thereby enhancing the clarity and reliability of the result. The data-driven nature of the Bayesian approach reveals the correlation and complex relationships between the parameters. To verify this, it is beneficial to obtain the correlation values between the parameters by defining the covariance matrix. The covariance matrix is calculated as follows:
@@
This equation provides a scaled representation highlighting the strength and direction of the linear relationships between variables independent of their units. Figure 4 visualizes these correlations for longitudinal and transverse electron and pion asymmetries, derived from the Monte Carlo minimization method and Bayesian analysis in the QWeak experiment.

### v2

@@ I Introduction
$A_{PV}=\frac{\frac{A_{\rm expt}}{P_{b}}-\sum_{i}f_{i}^{\rm bkgd}A_{i}^{\rm bkgd}}{1-\sum_{i}f_{i}^{\rm bkgd}},$ (1)
where $A_{\rm expt}$ is the experimentally measured asymmetry, which must be corrected for background processes characterized by fractional dilution factors $f_{i}^{\rm bkgd}$ and asymmetries $A_{i}^{\rm bkgd}$, and $P_{b}$ is the beam polarization.
@@
The basis of Bayesian statistics was first described in 1763 by Reverend Thomas Bayes and published by Richard Price on inverse probability [12]. In 1825, Pierre Simon Laplace published the Bayes’ theorem [13]. In the past 50 years, the ideas of inverse probability and Bayes’ theorem have been prominent tools in applied statistics although they are long-standing in mathematics. Bayesian methods of data analysis are now widely used across different fields of science such as ecology[14], social and behavioural sciences[15], genetics[16], medicine[17], educational research[18], epidemiology[19], organizational sciences[20],[21], modeling[22], nuclear physics[23], experimental data analysis[24], and the experimental particle physics[25]. However, Bayesian methods have not yet been applied in PVES experiments.
@@ II.1 Bayesian Analysis
The Bayesian approach studies conditional probability. When two occurring events A and B are dependent or conditional, the basic conditional probability can be written as
$P(B\cap A)=P(B|A)P(A),$ (2)
where ${P(B\cap A)}$ represents the probability of occurrence of both B and A, $P(B|A)$ is the probability of event B conditional on the occurrence of event A, and $P(A)$ is the probability of occurring event A. While, generically, $P(B|A)\not=P(A|B)$, however $P(B\cap A)=P(A\cap B)$, so we can also write
${P(A\cap B)}=P(A|B)P(B)$ (3)
which can be combined with 2, to yield
$P(A|B)=\frac{P(B|A)P(A)}{P(B)}.$ (4)
Equation 4 is Bayes’ rule. By extending these principles to the situation of data and model parameters, A would be the observation of specific data set $y$ and B is the realization of the model with parameters $\theta$. So, Equation 4 is written as follows:
$P(\theta|y)=\frac{P(y|\theta)P(\theta)}{P(y)}$ (5)
$P(\theta|y)$ is a conditional probability of the model parameters $\theta$ on the observation of data set $y$ and represents the posterior distribution. $P(y|\theta)$ is the conditional probability of the data given the model parameters and represents the likelihood function. Finally, $P(\theta)$ represents the probability of model parameter values in the population, known as the prior distribution and $P(y)$ is a normalizing factor. So, the posterior distribution is proportional to the likelihood function and the prior distribution. As a conclusion, the three basic steps of the typical Bayesian workflow are as follows:
  1. 1.
    Choosing the prior distribution, which is normally chosen before data collection, to represent available knowledge about a certain parameter in a statistical model.
  2. 2.
    Choosing the parameter information present in the observed data to determine the likelihood.
  3. 3.
    Combining the prior distribution and the likelihood function to form the posterior using Bayes’ theorem.
@@
This theorem demonstrates how beliefs are updated in light of new data and details how uncertainties regarding model parameters are quantified.
@@
To implement Bayesian analysis in the QWeak and MOLLER experiments, we quantify the prior distribution based on existing knowledge, use experimental data to form the likelihood function, and select a model to link these to the parameters of interest, the asymmetry components. Regarding the prior distribution, $P(\theta)$ in Equation 5, a noninformative prior knowledge is assumed for both experiments, represented by a Gaussian distribution with a very large uncertainty. This allows the measurements to be the dominant influence on the posterior distributions. Concerning the input data, $y$ in Equation 5, the measured asymmetry values are crucial. These values depend on the configuration of the experimental setup and the number of detectors, which are detailed in subsequent sections. Alongside the measured asymmetry values, other parameters, such as the beam polarization angle and spin variation, should also be considered as input data. Finally, the models that link the input data to the model parameters differ between the experimental setups.
The inherent complexities and computational challenges of working with these high-dimensional models necessitate the use of approximation techniques. These techniques aim to simplify reality, thereby enhancing our understanding of the system’s components. Various computational challenges in Bayesian inference have historically favoured frequentist methods until advancements in computational techniques allowed more efficient Bayesian methods. Techniques such as Approximate Bayesian Computation (ABC) [26], Integrated Nested Laplace Approximations (INLA) [27], and Variational Bayesian methods [28] have been developed. These methods generally approximate only marginal posterior distributions of individual parameters, not the joint distributions of multiple parameters with their potential correlations. MCMC methods, introduced by Gelfand and Smith in 1990 [29], have become prominent for their ability to approximate the joint posterior distribution through sampling. In these methods, which are computationally efficient and widely used in Bayesian analysis, the Monte Carlo part denotes the sampling process, and the Markov Chain part describes the mechanism for obtaining these samples.
@@ II.2 QWeak Experiment
In a shorter ancillary measurement of the QWeak experiment [33], the apparatus and experimental conditions were modified in two main ways compared to the weak-charge measurement: the beam energy was increased to 3.35 GeV to access the inelastic scattering kinematics of interest, and one of the main detectors was modified to enhance its sensitivity to pions. Due to the higher beam energy, a significant background was introduced in the main detectors, caused by negative pions produced in the target. Positively charged pions were swept out of the acceptance by the spectrometer’s magnetic field. Due to the high-rate integrating mode of the detector readout, it was not possible to separate the contributions of individual electrons and pions to the asymmetry measurement. Therefore, in the ancillary measurement, one of the main detectors was modified to enhance its sensitivity to pions in order to measure and correct for this pion background. The modification involved adding a 10.2 cm thick Pb absorber just upstream of the detector, which significantly attenuated the signal from scattered electrons without affecting the signal from the majority of pions. As a result, the asymmetry in MD7 was dominated by the incident pions, with a different mixture of electron and pion signals compared to the other seven main detectors.
@@
Here, $f_{\pi}^{i}$ represents the fractional yield of pions at MD $i$, and $A_{e(\pi)}^{L}$ and $A_{e(\pi)}^{T}$ denote the longitudinal and transverse asymmetries for electrons and pions, respectively. $\theta_{P}^{j}$ is the beam polarization angle for the run type $j$, and $f_{\text{NB}}^{i}$ is the neutral background yield fraction for MD $i$. The azimuthal angles, $\phi^{i}$, define the placement of the MDs with specific values like $\phi^{1}=0^{\circ}$ and $\phi^{2}=45^{\circ}$, among others. This model was first used in a Many-Worlds Monte Carlo minimization approach (a frequentist approach) to analyze the QWeak inelastic data, as implemented in [33] and [34], and will also be used to reanalyze the data using the Bayesian analysis method, incorporating all inputs from Ref. [33].
In the Many-Worlds Monte Carlo minimization approach, to extract the asymmetry components, $A_{e}^{L}$, $A_{e}^{T}$, $A_{\pi}^{L}$, and $A_{\pi}^{T}$, from the measured asymmetries in Equation 7, a value for each input quantity was randomly selected from a Gaussian distribution about their mean with widths equal to their uncertainties. These random values were then used to calculate the asymmetry in each MD $A_{i,j}^{\text{calc}}$ and for each polarization configuration via the following equation [34]:
$A_{i,j}^{\text{calc}}=(1-\tilde{f}_{\text{NB}}^{i})\biggl{[}$ $(1-\tilde{f}_{\pi}^{i})(A_{e}^{L}\cos\tilde{\theta}_{P}^{j}+A_{e}^{T}\sin\tilde{\theta}_{P}^{j}\sin\tilde{\phi}^{i})$ (8)
$+\tilde{f}_{\pi}^{i}(A_{\pi}^{L}\cos\tilde{\theta}_{P}^{j}+A_{\pi}^{T}\sin\tilde{\theta}_{P}^{j}\sin\tilde{\phi}^{i})\biggr{]},$
where a tilda over a quantity indicates a randomly selected value for that quantity. The function $\delta$ [34], where
$\delta_{\text{dof}}^{2}=\sum\left(A_{i,j}^{\text{meas}}-A_{i,j}^{\text{calc}}\right)^{2},$ (9)
was then minimized with respect to the asymmetries. This resulted in one possible set of values for each asymmetry, $A_{e}^{L}$, $A_{e}^{T}$, $A_{\pi}^{L}$, and $A_{\pi}^{T}$. The randomization and minimization process was repeated $10^{6}$ times, giving $10^{6}$ extracted values for each of the four asymmetries, which were used to shape the histograms shown in Fig. 1. The root mean squared of the resulting distributions was taken as their uncertainties. The same data will be reanalyzed using the Bayesian approach, and the corresponding results will be presented and discussed in the next section.
Figure 1: The distributions of the four asymmetries extracted through the Many-Worlds Monte Carlo minimization method, upper left: the asymmetry from electrons due to the longitudinal beam; upper right: the asymmetry from electrons due to the transverse beam; lower left: the asymmetry from pions due to the longitudinal beam; and lower right: the asymmetry from pions due to the transverse beam. It is important to note that there are differences in both sign and value between these histograms and the asymmetries presented in Table 1, due to the beam polarization sign and value not being corrected during the initial analysis. The correct signs and values are those listed in Table 1 [34].
@@ II.3 MOLLER Experiment
The experimental design differs significantly between the two; the QWeak data employed eight main detectors, with one acting as a pion detector, whereas the MOLLER experiment is planned to use a larger array of 84 main detector modules for ring 5 and 28 pion detector modules, necessitating an expanded Bayesian analysis framework. Note that in this paper, the analysis will focus on the main detector ring 5 and the pion detector. Eventually, we plan to expand this analysis to include all six rings of the main detectors, comprising 224 modules, the pion detector, and the showermax, each including 28 modules.
@@ III Validation of the Results
This section illustrates the results of implementing Bayesian analysis using the inputs and models from the QWeak and MOLLER experiments. For QWeak, validation is achieved by comparing these results with those obtained from the Many-Worlds Monte Carlo minimization approach [33, 34]. For MOLLER, validation is conducted by testing agreement with the input asymmetries in the simulations.
@@ III.1 Reanalysis of the QWeak Inelastic Experiment
Incorporating all inputs from the QWeak inelastic data, $A^{\text{meas}}$, $f_{\pi}$, $f_{\text{NB}}$, $\theta$, and $\phi$ (with $y$ as defined in Eq. 5), along with the model described by Eq. 7, the asymmetry values for Møller electrons and pions, $A_{e}^{L}$, $A_{e}^{T}$, $A_{\pi}^{L}$, and $A_{\pi}^{T}$ (with $\theta$ as defined in Eq. 5), have been calculated using the Many-Worlds Monte Carlo Minimization method [33, 34] and are now recalculated using Bayesian analysis.
@@
Figure 2: Comparative analysis of the measured inelastic QWeak and fitted asymmetry values across detectors for two data sets. On the left, the mixed data set is shown with measured asymmetry depicted as blue squares, compared with values derived through Bayesian inference (red circles) and Monte Carlo simulations (black triangles). On the right, the transverse data set is similarly represented. The data spans eight detectors in each plot, with error bars indicating the uncertainty for each data point. MD7 is highlighted with hollow symbols to indicate its distinct characteristics from the other MDs. Dashed lines in both data sets show the fitted sine wave in the Bayesian analysis, which is different from those obtained using the Monte Carlo method [33].
@@
This is explained by the mechanisms of both analysis methods. As outlined in the methodology, Bayes’ theorem involves an iterative process in which the posterior distribution from each iteration serves as the prior for the next, incorporating new data sets. This approach ensures that, in the absence of informative prior knowledge, the Bayesian analysis remains data-driven. Additionally, uncertainty is quantified by the width of the posterior distribution in the final step, which is shaped by the data and leads to more precise estimations.
In contrast, the Monte Carlo minimization method also explained in the methodology section, derives asymmetry components from measured data by randomly sampling input values from Gaussian distributions based on their means and uncertainties. These values are used to calculate asymmetries, and the method minimizes the squared deviations between measured and calculated asymmetries to estimate the unknown components. This process generates distributions for each component, as shown in Figure 1, and the RMS of these distributions is used to determine the uncertainty for each component, resulting in a tail in the histogram and larger uncertainty estimates due to the broader spread of possible values.
The results of this comparison suggest that assuming Gaussian distributions for inputs that are likely non-Gaussian, such as $f_{\pi}$ (which is strictly bound in the interval $[0,1]$), is problematic in the Monte Carlo minimization method. In a test study, when we reduced the uncertainty on $f_{\pi}$ by one order of magnitude, the tails of the distributions in Figure 1 disappeared, and the uncertainties became closer to those obtained from the Bayesian analysis. This study confirmed that the assumption of Gaussianity for all inputs is imprecise. Furthermore, the method used to calculate uncertainty was found to be inaccurate. The RMS of the distributions of squared deviations between measured and calculated asymmetries is not a precise way to quantify uncertainty. Addressing these issues should ultimately lead to consistent results from both methods.
@@ III.2 Analysis of the MOLLER Experiment
Incorporating the simulations and assumptions from the MOLLER experiment, the input data include $A^{\text{mock}}$, $f^{i}_{\pi}$, $\theta_{P}^{j}$, $\phi_{P}^{j}$, $A_{ij}$, $B_{ij}$, $C_{ij}$, and $D_{ij}$ (i.e. these are the $y$ in Eq. 5). The parameters to be extracted are the asymmetry values $A_{e}^{L}$, $A_{e}^{T}$, $A_{\pi}^{L}$, and $A_{\pi}^{T}$ (i.e. these are the $\theta$ in Eq. 5). The model, specified by Eq. 11, is applied separately to the pion and main detectors for both longitudinal and transverse measurements. Consequently, the asymmetry values are extracted using four equations: two for the pion detector and two for the main detector, covering both longitudinal and transverse measurements. The results of the extracted asymmetry values and their associated uncertainties in ppb are summarized in Table 2 and compared with the inputs of the analysis.
@@
Another way to verify the results is by substituting the asymmetry values along with other parameters into the right-hand side of Eq. 11, computing and deriving another set of asymmetry values, referred to as the fitted asymmetry values, similar to the case of Qweak. Figures 3 and 4 provide a comparative analysis of fitted versus mock asymmetry for the main and pion detectors in both the longitudinal and transverse data sets. In the top plots, the blue squares represent the average of mock asymmetry values, while the red circles denote the average of fitted asymmetry values obtained through Bayesian analysis. Due to the large scale of the asymmetry values, the error bars are too small to be visible in some of these plots. To address this, the middle plots subtract the mock asymmetry average from both the mock and fitted asymmetry values, effectively re-centering the data around zero. This re-centering provides a clearer visualization of the error bars, illustrating the measurement precision. In these plots, the blue dots represent the residual mock asymmetry values, which are zero as they are subtracted from themselves. The red circles display the residual fitted asymmetry values, with the error bars indicating the uncertainties. The bottom plots present histograms of the normalized residuals, defined as the difference between the fitted and mock asymmetry values, normalized by the uncertainty in the mock asymmetry values. These histograms help assess the normality and spread of these residuals.
@@ IV Discussion
Understanding the correlation between parameters is essential for revealing the complex relationships that govern the data. By quantifying how different variables co-vary, we can gain deeper insights into their interdependencies and improve the precision of parameter estimates. To achieve this, the covariance matrix provides a valuable tool for evaluating these correlations. The covariance matrix is calculated as follows:
@@
This equation provides a scaled representation highlighting the strength and direction of the linear relationships between variables independent of their units.
Figure 5 visualizes these correlations for longitudinal and transverse electron and pion asymmetries, derived from the Monte Carlo minimization method and Bayesian analysis in the QWeak experiment.